# Very Massive Stars and High N/O: A Tale of the Nitrogen-enriched Super Star Cluster in the Sunburst Arc

### v1

@@ 4.1.2 Chemical enrichment
We find that these nitrogen-enriched gas cells undergo two dynamic phases. They first sink to the center of the central massive cluster, as the separation $|\mathbf{x}_{\rm gas}-\mathbf{x}_{\rm center}|$ plummets to $\sim 0.1\,\rm pc$ for $t-t_{0}\sim 3-4\,\rm Myr$ (where they also get well mixed with stellar wind material). Meanwhile, the typical pressure of these gas cells rises to $\gtrsim 10^{9}\,\rm K\,cm^{-3}$ and the typical density reaches $\gtrsim 10^{6}\,m_{\rm p}/\rm cm^{3}$. As expected, they are rapidly enriched by stellar winds at $\sim 3\,\rm Myr$ (SN explosions are yet to occur at this young age). At this point, the velocity dispersion of these gas cells is $\sim 10-20\,\rm km/s$, smaller than but on the same order of magnitude as the velocity dispersion of the LyC cluster as inferred from the width of nebular lines ($\sim 40\,\rm km/s$, Vanzella et al., 2022). This also implies that these gas cells are gravitationally bound to the cluster. The sound speed stays low $\sim 1\,\rm km/s$ for $\sim 3\,\rm Myr$, but then rapidly rises to $\sim 10\,\rm km/s$ afterwards, which implies a photo-ionization process. The time evolution of the density and pressure for these gas cells reveals a nearly isothermal gravitational collapse followed by photo-heating due to the stars.
@@
After nitrogen enrichment, the high-pressure gas cells are quickly “kicked out” from the cluster’s center to $\gtrsim 100\,\rm pc$ away. The gas is no longer bound to the cluster as the velocity dispersion surges to $\sim 200\,\rm km/s$. Losing external pressure confinement, the pressure and density of these gas cells drop dramatically, while the sound speed $\sim 10\,\rm km/s$ is consistent with photo-ionization as noted above. This represents the process where stellar feedback efficiently drives away the polluted gas (natal gas mixed with stellar wind material). During this phase, the nitrogen-enriched gas still maintains a high N/O ratio, but with a large variance that is potentially caused by diffusive mixing of metals.
@@
In summary, the nitrogen-enrichment appears as a temporary phase of star cluster formation from GMC collapse. In terms of mass, the nitrogen-enriched gas is a small subset of the gas complex (also see discussion in § 4.1.3). The enriched gas is typically dense ($\gtrsim 10^{5}\,\rm cm^{-3}$) and highly-pressurized ($\gtrsim 10^{9}\,\rm K\,cm^{-3}$), in agreement with what has been observed in the Sunburst LyC cluster and other high-redshift nitrogen-enriched galaxies (Pascale & Dai, 2024; Ji et al., 2024; Topping et al., 2024; Yanagisawa et al., 2024). This gas originates from the natal GMC and settles down at the cluster center following a nearly isothermal gravitational collapse, but is then heated by stellar feedback and eventually ejected from the cluster.
@@
Recent studies (Rizzuti et al., 2025; McClymont et al., 2025) suggest that elevated N/O ratio is a temporary phenomenon in the ISM of high-redshift galaxies, and arises from differential winds. Unlike the nitrogen-enrichment process we investigate here, the relevant spatial scale is much larger than the proximity of newborn super star clusters, and the relevant timescale more aligns with the longer evolution of AGB stars rather than with the rapid evolution of VMSs. It is unclear if that picture is compatible with the intriguing observation that the nitrogen-enriched gas is almost always as dense as $\gtrsim 10^{5}\,\rm cm^{-3}$ in observed galaxies and clumps (references listed above).
@@
4.1.3 Nitrogen mass budget
@@
In Fig. 9, we track the mass evolution of the high-pressure ($10^{9}\,\rm K\,cm^{-3}$) gas $m_{\rm highP}$, which is observed to be nitrogen-enriched (Pascale et al., 2023, and also found in our simulation). For three fiducial simulations at $Z=0.2\,Z_{\odot}$, $m_{\rm highP}$ can reach $\sim 10^{5}\,M_{\odot}$ at $t-t_{0}\sim 2\,\rm Myr$ (though this peak value slightly decreases for less compact GMCs), while it remains $m_{\rm highP}\sim 10^{4}\,M_{\odot}$ at $t-t_{0}\sim 4\,\rm Myr$ (i.e., the peak time of N/O anomaly, see Fig. 6). However, the high-pressure gas is removed after $t-t_{0}\sim 4\,\rm Myr$.
@@
In summary, we find that the mass of the high-pressure gas is $\sim 10^{4}\,M_{\odot}$ at the culmination of nitrogen enrichment in the simulation, which is a tiny fraction of the total GMC mass ($3\times 10^{7}\,M_{\odot}$). Despite its insignificance in terms of mass, this gas is in very close proximity to massive stars at the center of the cluster, such that it may absorb a sizable or even dominant fraction of the stellar ionizing radiation, and hence dominate the observed nebular emission from the high-ionization zone. The two-component photo-ionization model in Pascale et al. (2023) captures this. The high-pressure gas is a mixture of natal gas and, predominantly, stellar wind ejecta, such that most nitrogen there comes from nitrogen-rich winds of massive stars (also see Fig. 2), particularly of the VMSs. While the high-pressure gas is indeed nitrogen-enriched, the majority of the gas in the GMC is “low-pressure” and lacks a dramatic N/O elevation, as also reflected in Fig. 5. Interestingly, Ji et al. (2024) found that the low-density gas of GS_3073 is not enriched, similar to the prediction of the scenario here.
4.2 Simulations with different GMC masses
@@
Fig. 10 shows the time evolution of star formation and the chemical abundances of the high-pressure gas for these control simulations (with the same convention as Fig. 6). The final stellar mass reaches $\sim 5\times 10^{7}\,M_{\odot}$ for GMCs of $10^{8}\,M_{\odot}$ in contrast to $\sim 2\times 10^{6}\,M_{\odot}$ for GMCs of $10^{7}\,M_{\odot}$, showing that the star formation efficiency (SFE) increases (from $\sim 20\%$ to $\sim 50\%$) with the GMC mass, and more fundamentally, the GMC initial mean surface density as found by previous studies (e.g., Grudić et al., 2018; Chevance et al., 2023, and reference therein).
@@
The simulation M1e8_R60_Z0.2 with the highest surface density for the natal GMC also has the smallest UV half-light radius ($\sim 2-3\,\rm pc$), which is close to the half-mass radius of the central cluster. We also visualize the stellar distribution of this simulation in the lower panels of Fig. 7, where almost all stars belong to the central massive cluster (though there are indeed many smaller satellite clusters). All other simulations typically resemble the upper panels of Fig. 7, where sub-clusters are gravitationally unbound from each other and scatter throughout the entire simulation domain (so the UV half-light radius of the system is much larger than the half-mass radii of sub-clusters). The simulation M1e8_R60_Z0.2 is an example of hierarchical mergers where sub-clusters are gravitationally bound to the central cluster that they will eventually merge into (Shi et al., 2024).
We further check the chemical evolution of the high-pressure ($10^{9}\,\rm K\,cm^{-3}$) gas in these two sets of simulations. Simulations with $M_{\rm cloud}=10^{8}\,M_{\odot}$ have similar N/O and C/O signals to those of the fiducial simulations, but enrichment occurs earlier, in only $\sim 2-3\,\rm Myr$. However, the high-pressure gas in simulations with $M_{\rm cloud}=10^{7}\,M_{\odot}$ can only rise to $\log(\rm N/O)\sim-1$ throughout the evolution, well below the N/O value found in Pascale et al. (2023). As a result, these less massive and less compact GMCs are unlikely to be the progenitor of the LyC cluster. In fact, less compact GMCs may not even confine gas with $p/k_{\rm B}>10^{9}\,{\rm K}\,{\rm cm}^{-3}$ in its star-forming complex. A simple, order-of-magnitude argument is that the self-gravity, $g\sim GM_{\rm cloud}/R_{\rm cloud}^{2}\sim\pi G\Sigma$, should be greater than the buoyancy, $\nabla p/\rho\sim(p/R_{\rm cloud})/(3M_{\rm cloud}/4\pi R_{\rm cloud}^{3})\sim p/(3\Sigma/4)$. Insert $p=k_{\rm B}\cdot 10^{9}\,{\rm K}\,{\rm cm}^{-3}$, we find $\Sigma\gtrsim(4p/3\pi G)^{1/2}\gtrsim 1\,{\rm g}\,{\rm cm}^{-2}$, roughly corresponds to the clouds we are simulating.
Figure 11: Simulations without VMSs (i.e., FIRE-only). Conventions are the same as Fig. 6. Top row - FIRE simulations but consider the “failed” SNe (corresponding to ZAMS mass of $\sim 25-70\,M_{\odot}$, see § 3.3). Bottom row - simulations with default FIRE setups.
4.3 Simulations without VMSs
@@
In these simulations, chemical feedback is completely determined by the sub-grid models of FIRE-3 (Hopkins et al., 2023). We run two versions of FIRE-only simulations. First, we consider “failed SNe” such that there is no SN feedback before $t_{\rm CC}\sim 8\,\rm Myr$. This is also the modification that we made for our FIRE+VMS model described earlier (§ 3.3), but the difference is that no VMSs are modeled and the IMF extends only to $100\,M_{\odot}$ (so PPISNe at $\sim 3\,\rm Myr$ are also missing). This leads to weaker stellar feedback. This is reflected in the star formation history, where the stellar mass accumulates to $\sim 2\times 10^{7}\,M_{\odot}$, roughly twice as much compared with the FIRE+VMS runs (§ 4.1).
@@
This is consistent with the theory (Murray et al., 2010; Fall et al., 2010; Matzner & Jumper, 2015) and simulations (Grudić et al., 2018; Chevance et al., 2023) that the SFE is determined by the competition between self-gravity, $F_{\rm g}\sim G\,M_{\rm cloud}\,M_{\rm gas}/R_{\rm cloud}^{2}\sim\pi\,G\,M_{\rm gas}\,\Sigma_{\rm cl}$, and the feedback force, characterized by $F_{\rm fb}\sim(\dot{p}/m)M_{\star}$, where $\dot{p}/m$ is the momentum ejection rate per unit mass from stars. By removing the VMS populations, $\dot{p}/m$ for the stellar population decreases since VMSs have a higher L/M ratio, which leads to the dominance of self-gravity and a higher SFE in the new balance. The dominance of self-gravity also explains the UV half-light radii of the FIRE-only runs, which are $\sim 2-3\,\rm pc$ after $t-t_{0}\sim 4\,\rm Myr$, significantly smaller than in simulations with VMSs (Fig. 6).
@@
Another version of the FIRE-only runs keeps all the original setups of FIRE-3. We find no significant difference in terms of star formation when compared to the FIRE-only runs with “failed SNe.” SN feedback takes effect when the age of the stellar population (i.e., $t-t_{0}$) reaches $3.7\,\rm Myr$ (Hopkins et al., 2023), too late to impact star formation.
@@
The differences between the two simulation versions are more significant in chemical feedback. With FSN implemented, chemical feedback before $t_{\rm cc}\sim 8\,\rm Myr$ in the FIRE-only runs is dominated by stellar winds from stars with $100\,M_{\odot}$. We indeed see nitrogen elevation to $\log(\rm N/O)\gtrsim-0.5$ at $t-t_{0}\sim 5\,\rm Myr$, compatible with our toy model compiled with PARSEC data (Fig. 2), but the timing of this nitrogen enrichment event $\sim 5\,\rm Myr$ is later than the observed $2-4\,\rm Myr$ range (Pascale et al., 2023). Moreover, the C/O ratio decreases with time, down to $\sim-1.4$ at $t-t_{0}\sim 5\,\rm Myr$, which becomes inconsistent with observations. Like in the fiducial simulations (Fig. 6), high-pressure gas is removed after nitrogen enrichment, but this takes place later at $t-t_{0}\sim 6\,\rm Myr$. Still, this is before the onset of CCSNe in this simulation, ruling out the possibility that the high-pressure gas is removed by SN feedback.
@@
The default FIRE-3 simulations do not reproduce the chemical feedback either. Although there is nitrogen enrichment at $t-t_{0}\sim 5\,\rm Myr$, it does not reach the observed range. Given the difference between the previous simulations with FSN, the nitrogen enrichment should be “interrupted” by oxygen yield from SN explosions. The C/O ratio here, however, surges to $\log({\rm C/O})>0$, which is also consistent with observations. This specific result is related to the SN yield model of FIRE-3 (Hopkins et al., 2023), where a high abundance of carbon is produced in early SN explosions.
@@
4.4 VMS feedback tests
Fig. 8 shows that the soon-to-be nitrogen-enriched gas in the simulation first sinks into the central massive cluster but is quickly ejected. We now investigate the exact mechanism (radiation, winds, or SNe) to eject (and pressurize) this gas. To that end, we run simulations in which certain feedback physics is disabled to determine the exact reason. This is implemented by multiplying the associated energy by a tiny factor of $10^{-10}$ (a factor of $10^{-5}$ for the wind momentum), while the mass outflow rate and related chemical feedback are untouched. Here we run three versions of simulations that 1) disable kinetic feedback from VMS SNe by setting $E_{\rm SN}\approx 0$ for those; 2) disable kinetic feedback from both SNe and winds from VMSs, by setting $E_{\rm SN}\approx E_{\rm wind}\approx 0$; 3) further disable all radiative feedback from stars (both FIRE stellar populations and VMSs), such that there is no stellar feedback before $\sim 8\,\rm Myr$ except for the rather weak stellar winds from FIRE stellar populations. Following the convention of Fig. 6, the results are shown in Fig. 12.

### v2

@@ IV.1.2 Chemical enrichment
We find that the nitrogen-enriched gas undergoes two dynamical phases. Initially, it sinks toward the center of the massive cluster as the separation $|\mathbf{x}_{\rm gas}-\mathbf{x}_{\rm center}|$ drops to $\sim 0.1\,\rm pc$ for $t-t_{0}\!\sim\!3$$4\,\rm Myr$, where it mixes with stellar-wind material. This agrees with Rivera-ThorsenChisholmWelch_2024A&A...690A.269R that the enriched gas should be condensed inside the cluster and close to the most massive stars. During this stage, the pressure rises to $\gtrsim 10^{9}\,\rm K\,cm^{-3}$ and the density to $\gtrsim 10^{6}\,m_{\rm p}\,{\rm cm^{-3}}$. Enrichment occurs at $\sim 3\,\rm Myr$, before SNe, with velocity dispersion $\sim 10$$20\,\rm km\,s^{-1}$—comparable to but below that inferred for the LyC cluster from nebular-line widths ($\sim 40\,\rm km\,s^{-1}$(VanzellaCastellanoBergamini_2022A&A...659A...2V), indicating the gas remains gravitationally bound. The sound speed remains $\sim 1\,\rm km\,s^{-1}$ until $\sim 3\,\rm Myr$, then rises to $\sim 10\,\rm km\,s^{-1}$, signaling photoionization. The evolution of density and pressure thus reflects an isothermal collapse followed by stellar photoheating.
@@
After enrichment, the high-pressure gas is expelled to $\gtrsim 100\,\rm pc$, with velocity dispersion increasing to $\sim 200\,\rm km\,s^{-1}$, unbinding it from the cluster. As external confinement is lost, both pressure and density decline, while the $\sim 10\,\rm km\,s^{-1}$ sound speed remains consistent with photoionized gas. This phase marks the efficient removal of polluted gas (natal material mixed with stellar winds). The nitrogen-enriched gas retains a high N/O ratio but with a large variance, likely due to turbulent mixing of metals.
@@
In summary, the nitrogen-enrichment appears as a temporary phase of star cluster formation from GMC collapse. In terms of mass, the nitrogen-enriched gas is a small subset of the gas complex (also see discussion in § IV.1.3). The enriched gas is typically dense ($\gtrsim 10^{5}\,\rm cm^{-3}$) and highly-pressurized ($\gtrsim 10^{9}\,\rm K\,cm^{-3}$), in agreement with what has been observed in the Sunburst LyC cluster and other high-redshift nitrogen-enriched galaxies (PascaleDai_2024ApJ...976..166P; JiUblerMaiolino_2024MNRAS.535..881J; ToppingStarkSenchyna_2024MNRAS.529.3301T; YanagisawaOuchiWatanabe_2024ApJ...974..266Y). This gas originates from the natal GMC and settles down at the cluster center following a nearly isothermal gravitational collapse, but is then heated by stellar feedback and eventually ejected from the cluster.
@@
Recent studies (RizzutiMatteucciMolaro_2025A&A...697A..96R; McClymontTacchellaSmith_2025arXiv250708787M) suggest that elevated N/O ratio is a temporary phenomenon in the ISM of high-redshift galaxies, and arises from differential winds. Unlike the nitrogen-enrichment process we investigate here, the relevant spatial scale is much larger than the proximity of newborn super star clusters, and the relevant timescale more aligns with the longer evolution of AGB stars rather than with the rapid evolution of VMSs. It is unclear if that picture is compatible with the intriguing observation that the nitrogen-enriched gas is almost always as dense as $\gtrsim 10^{5}\,\rm cm^{-3}$ in observed galaxies and clumps (references listed above).
@@
IV.1.3 Nitrogen mass budget
@@
In Fig. 10, we track the mass evolution of the high-pressure ($10^{9}\,\rm K\,cm^{-3}$) gas $m_{\rm highP}$, which is observed to be nitrogen-enriched (PascaleDaiMcKee_2023ApJ...957...77P, and also found in our simulation). For three fiducial simulations at $Z=0.2\,Z_{\odot}$, $m_{\rm highP}$ can reach $\sim 10^{5}\,M_{\odot}$ at $t-t_{0}\sim 2\,\rm Myr$ (though this peak value slightly decreases for less compact GMCs), while it remains $m_{\rm highP}\sim 10^{4}\,M_{\odot}$ at $t-t_{0}\sim 4\,\rm Myr$ (i.e., the peak time of N/O anomaly, see Fig. 7). However, the high-pressure gas is removed after $t-t_{0}\sim 4\,\rm Myr$.
@@
In summary, we find that the mass of the high-pressure gas is $\sim 10^{4}\,M_{\odot}$ at the culmination of nitrogen enrichment in the simulation, which is a tiny fraction of the total GMC mass ($3\times 10^{7}\,M_{\odot}$). Despite its insignificance in terms of mass, this gas is in very close proximity to massive stars at the center of the cluster, such that it may absorb a sizable or even dominant fraction of the stellar ionizing radiation, and hence dominate the observed nebular emission from the high-ionization zone. The two-component photo-ionization model in PascaleDaiMcKee_2023ApJ...957...77P captures this. The high-pressure gas is a mixture of natal gas and, predominantly, stellar wind ejecta, such that most nitrogen there comes from nitrogen-rich winds of massive stars (also see Fig. 3), particularly of the VMSs. While the high-pressure gas is indeed nitrogen-enriched, the majority of the gas in the GMC is “low-pressure” and lacks a dramatic N/O elevation, as also reflected in Fig. 6. Interestingly, JiUblerMaiolino_2024MNRAS.535..881J found that the low-density gas of GS_3073 is not enriched, similar to the prediction of the scenario here.
IV.2 Simulations with different GMC masses
@@
Fig. 11 shows the time evolution of star formation and the chemical abundances of the high-pressure gas for these control simulations (with the same convention as Fig. 7). The final stellar mass reaches $\sim 5\times 10^{7}\,M_{\odot}$ for GMCs of $10^{8}\,M_{\odot}$ in contrast to $\sim 2\times 10^{6}\,M_{\odot}$ for GMCs of $10^{7}\,M_{\odot}$, showing that the star formation efficiency (SFE) increases (from $\sim 20\%$ to $\sim 50\%$) with the GMC mass, and more fundamentally, the GMC initial mean surface density as found by previous studies (e.g., GrudicHopkinsFaucher-Giguere_2018MNRAS.475.3511G; ChevanceKrumholzMcLeod_2023ASPC..534....1C, and reference therein).
@@
The simulation M1e8_R60_Z0.2 with the highest surface density for the natal GMC also has the smallest UV half-light radius ($\sim 2-3\,\rm pc$), which is close to the half-mass radius of the central cluster. We also visualize the stellar distribution of this simulation in the lower panels of Fig. 8, where almost all stars belong to the central massive cluster (though there are indeed many smaller satellite clusters). All other simulations typically resemble the upper panels of Fig. 8, where sub-clusters are gravitationally unbound from each other and scatter throughout the entire simulation domain (so the UV half-light radius of the system is much larger than the half-mass radii of sub-clusters). The simulation M1e8_R60_Z0.2 is an example of hierarchical mergers where sub-clusters are gravitationally bound to the central cluster that they will eventually merge into (ShiKremerHopkins_2024ApJ...969L..31S).
We further check the chemical evolution of the high-pressure ($10^{9}\,\rm K\,cm^{-3}$) gas in these two sets of simulations. Simulations with $M_{\rm cloud}=10^{8}\,M_{\odot}$ have similar N/O and C/O signals to those of the fiducial simulations, but enrichment occurs earlier, in only $\sim 2-3\,\rm Myr$. However, the high-pressure gas in simulations with $M_{\rm cloud}=10^{7}\,M_{\odot}$ can only rise to $\log(\rm N/O)\sim-1$ throughout the evolution, well below the N/O value found in PascaleDaiMcKee_2023ApJ...957...77P. As a result, these less massive and less compact GMCs are unlikely to be the progenitor of the LyC cluster. In fact, less compact GMCs may not even confine gas with $p/k_{\rm B}>10^{9}\,{\rm K}\,{\rm cm}^{-3}$ in its star-forming complex. A simple, order-of-magnitude argument is that the self-gravity, $g\sim GM_{\rm cloud}/R_{\rm cloud}^{2}\sim\pi G\Sigma$, should be greater than the buoyancy, $\nabla p/\rho\sim(p/R_{\rm cloud})/(3M_{\rm cloud}/4\pi R_{\rm cloud}^{3})\sim p/(3\Sigma/4)$. Insert $p=k_{\rm B}\cdot 10^{9}\,{\rm K}\,{\rm cm}^{-3}$, we find $\Sigma\gtrsim(4p/3\pi G)^{1/2}\gtrsim 1\,{\rm g}\,{\rm cm}^{-2}$, roughly corresponds to the clouds we are simulating.
Figure 12: Simulations without VMSs (i.e., FIRE-only). Conventions are the same as Fig. 7. Top row - FIRE simulations but consider the “failed” SNe (corresponding to ZAMS mass of $\sim 25-70\,M_{\odot}$, see § III.3). Bottom row - simulations with default FIRE setups.
IV.3 Simulations without VMSs
@@
In these simulations, chemical feedback is completely determined by the sub-grid models of FIRE-3 (HopkinsWetzelWheeler_2023MNRAS.519.3154H). We run two versions of FIRE-only simulations. First, we consider “failed SNe” such that there is no SN feedback before $t_{\rm CC}\sim 8\,\rm Myr$. This is also the modification that we made for our FIRE+VMS model described earlier (§ III.3), but the difference is that no VMSs are modeled and the IMF extends only to $100\,M_{\odot}$ (so PPISNe at $\sim 3\,\rm Myr$ are also missing). This leads to weaker stellar feedback. This is reflected in the star formation history, where the stellar mass accumulates to $\sim 2\times 10^{7}\,M_{\odot}$, roughly twice as much compared with the FIRE+VMS runs (§ IV.1).
@@
This is consistent with the theory (MurrayQuataertThompson_2010ApJ...709..191M; FallKrumholzMatzner_2010ApJ...710L.142F; MatznerJumper_2015ApJ...815...68M) and simulations (GrudicHopkinsFaucher-Giguere_2018MNRAS.475.3511G; ChevanceKrumholzMcLeod_2023ASPC..534....1C) that the SFE is determined by the competition between self-gravity, $F_{\rm g}\sim G\,M_{\rm cloud}\,M_{\rm gas}/R_{\rm cloud}^{2}\sim\pi\,G\,M_{\rm gas}\,\Sigma_{\rm cl}$, and the feedback force, characterized by $F_{\rm fb}\sim(\dot{p}/m)M_{\star}$, where $\dot{p}/m$ is the momentum ejection rate per unit mass from stars. By removing the VMS populations, $\dot{p}/m$ for the stellar population decreases since VMSs have a higher L/M ratio, which leads to the dominance of self-gravity and a higher SFE in the new balance. The dominance of self-gravity also explains the UV half-light radii of the FIRE-only runs, which are $\sim 2-3\,\rm pc$ after $t-t_{0}\sim 4\,\rm Myr$, significantly smaller than in simulations with VMSs (Fig. 7).
@@
Another version of the FIRE-only runs keeps all the original setups of FIRE-3. We find no significant difference in terms of star formation when compared to the FIRE-only runs with “failed SNe.” SN feedback takes effect when the age of the stellar population (i.e., $t-t_{0}$) reaches $3.7\,\rm Myr$ (HopkinsWetzelWheeler_2023MNRAS.519.3154H), too late to impact star formation.
@@
The differences between the two simulation versions are more significant in chemical feedback. With FSN implemented, chemical feedback before $t_{\rm cc}\sim 8\,\rm Myr$ in the FIRE-only runs is dominated by stellar winds from stars with $100\,M_{\odot}$. We indeed see nitrogen elevation to $\log(\rm N/O)\gtrsim-0.5$ at $t-t_{0}\sim 5\,\rm Myr$, compatible with our toy model compiled with PARSEC data (Fig. 3), but the timing of this nitrogen enrichment event $\sim 5\,\rm Myr$ is later than the observed $2-4\,\rm Myr$ range (PascaleDaiMcKee_2023ApJ...957...77P). Moreover, the C/O ratio decreases with time, down to $\sim-1.4$ at $t-t_{0}\sim 5\,\rm Myr$, which becomes inconsistent with observations. Like in the fiducial simulations (Fig. 7), high-pressure gas is removed after nitrogen enrichment, but this takes place later at $t-t_{0}\sim 6\,\rm Myr$. Still, this is before the onset of CCSNe in this simulation, ruling out the possibility that the high-pressure gas is removed by SN feedback.
@@
The default FIRE-3 simulations do not reproduce the chemical feedback either. Although there is nitrogen enrichment at $t-t_{0}\sim 5\,\rm Myr$, it does not reach the observed range. Given the difference between the previous simulations with FSN, the nitrogen enrichment should be “interrupted” by oxygen yield from SN explosions. The C/O ratio here, however, surges to $\log({\rm C/O})>0$, which is also consistent with observations. This specific result is related to the SN yield model of FIRE-3 (HopkinsWetzelWheeler_2023MNRAS.519.3154H), where a high abundance of carbon is produced in early SN explosions.
@@
IV.4 VMS feedback tests
Figure 9 shows that nitrogen-enriched gas in the simulation rapidly sinks into the central cluster before being expelled. To identify the feedback mechanism responsible (radiation, winds, or SNe), we rerun the simulations with specific feedback channels disabled. This is achieved by reducing the associated energy by a factor of $10^{-10}$ ($10^{-5}$ for wind momentum) while keeping the mass outflow and chemical feedback unchanged. We perform three tests: (1) VMS SNe with kinetic feedback suppressed ($E_{\rm SN}\approx 0$); (2) both VMS SNe and wind kinetic feedback disabled ($E_{\rm SN}\approx E_{\rm wind}\approx 0$); and (3) all stellar radiative feedback turned off (for both FIRE SSPs and VMSs), leaving only the weak winds from FIRE populations until $\sim 8\,\rm Myr$. Results for these runs are shown in Fig. 13, following the format of Fig. 7.